# Simulation-based inference of galaxy properties from JWST pixels

### v1

@@ 3.2 Amortised neural inference
We used a neural density estimator known as NFs to estimate the posterior distributions for the stellar population properties. This method transforms variables described by a simple base distribution, such as a multivariate Gaussian, through a series of parametrised invertible transformations. The parameters of these transformations are trained to minimize the Kullback–Leibler divergence between the target distribution and the model’s estimate. For a more detailed explanation, see Jimenez Rezende & Mohamed (2015) and Hahn & Melchior (2022).
We used a particular implementation of NFs known as MAF (masked autoregressive flow; Papamakarios et al., 2017), implemented with the SBI module (Tejero-Cantero et al., 2020). Following the example of Hahn & Melchior (2022), we selected 15 MADE blocks (masked autoencoder for distribution estimation), each containing two hidden layers with 500 units. We split the data into training and validation sets in a 90/10 ratio, and used the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $5\cdot 10^{-4}$. To prevent overfitting, we monitored the model’s performance on the validation set, and halted the training if the validation likelihood did not improve after $20$ epochs.
We performed two trainings for the two different simulations, with parametric (a) and non-parametric (b) SFHs. The output of the network are the posterior distributions for the following properties: the formed stellar mass , the surviving stellar mass including remnants, the average SFR for the last $100$ Myr, the metallicity $[M/H]$ and the dust attenuation index $A_{V}$. For the SFHs, in (a) we inferred $\tau$ and $t_{i}$, both in Gyr, while in (b) we inferred the normalised time $(t_{x}/t_{\rm{universe(z)}})$ at which $25\%$, $50\%$ and $75\%$ of the total stellar mass were formed.
We discuss the possibility of also retrieving the redshift pixel-by-pixel, as done in (Pérez-González2023) in Sect.7. However, in this work we introduced the already mentioned integrated spectroscopic redshifts as an input in the network together with the photometry.
The training was performed with $1.000.000$ simulated CSPs. We generated another $10.000$ for testing purposes.
@@ 3.3 Fitting of JADES galaxies
We use the trained model to fit the galaxies introduced in Sect. 2. We compute the noise of the galaxies in each pixel and filter as $\sqrt{\sigma^{2}+\sigma_{b}^{2}}$, where $\sigma$ is the flux uncertainty from the mosaics and $\sigma_{b}$ the standard deviation of the background signal. We refer to the NIRCam IDs of the galaxies throughout the paper. All spectroscopic redshifts are extracted from catalogues (Dahlen et al., 2010; Guo et al., 2012; Oesch et al., 2023; Bunker et al., 2024). We input into the network the photometry with uncertainties (in AB magnitudes), repeating again the cut at a $1\sigma$ depth limit for each filter as done for the simulations, as well as the redshifts. We thus obtain posterior distributions for each pixel for (a) $\log_{10}(M_{*}/\rm{M}_{\odot})$, $\log_{10}(M_{*}^{\rm{formed}}/\rm{M}_{\odot})$, $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)), $\tau$, $t_{i}$, [$M$/H] and $A_{V}$; and (b) $\log_{10}(M_{*}/\rm{M}_{\odot})$, $\log_{10}(M_{*}^{\rm{formed}}/\rm{M}_{\odot})$, $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)), $t_{25\%}$, $t_{50\%}$, $t_{75\%}$, [$M$/H] and $A_{V}$.
The results are organised as follows. In Sect. 4 we analyse the performance of the model in the simulated test set. In Sect. 5 we fit individual pixels from galaxies of the JADES survey. We then inspect the posterior distributions at different redshifts (Sect. 5.1), comparing them with other SED fitting codes. Then, in Sect. 6, we obtain maps for stellar population properties of selected galaxies. Finally, in Sect. 6.2 we perform a statistical study and compare the integrated stellar masses with their pixel-based analogues to investigate the outshining phenomenon.
@@ 4 Recovering properties of the simulated stellar populations
We train the neural density estimator during $\sim 1$ hour and $30$ min. We obtain posterior probability distributions for each stellar population parameter for the full test set, from the photometry with the associated uncertainties. It takes $\sim 0.04$ s to obtain $500$ samples of the distributions for all the properties for one simulation. All the time computations were performed on a single core of an Apple M3 Max CPU, part of an ARM-based 14-core system with 36 GB of unified memory. However, this model natively supports GPU execution, which can further reduce inference time.
@@
To assess the accuracy of the model, we compute the medians of the posteriors inferred for the parameters and compare them with the true simulated values for the full test sample with S/N averaged in the filters F277W, F356W and F [^4] W ($\overline{\rm{S/N}}$ from now on) higher than 1. This ensures we do not show simulations that are completely above of the observational capabilities of ACS and NIRCam, with upper limits for most of the filters. In Table 3 we include the $R^{2}$ values [^3] Also known as coefficient of determination $R^{2}=1-\frac{\sum_{i=1}^{n}\left(\theta_{i}-\widehat{\theta}_{i}\right)^{2}}{\sum_{i=1}^{n}\left(\theta_{i}-\overline{\theta}_{i}\right)^{2}}$, not to be confused with the square of the Pearson’s coefficient. for the training using parametric (a) and non-parametric (b) SFHs. The performance is relatively good for the stellar mass and the dust attenuation index, but worse for the SFR, metallicity and ages, reproducing the degeneracies we know exist when inferring these properties from photometry. As expected, the $R^{2}$ values for the SFR and [$M$/H] are higher for parametric SFHs. However, while $t_{i}$ can be well constrained, the accuracy of the medians of the posterior distribution for $\tau$ are extremely low. This can be explained not only because it is difficult to retrieve from photometry, but also because it is very degenerated with $t_{i}$ (Conroy, 2013). Instead, we find that by combining $\tau$ and $t_{i}$ samples to obtain mass-weighted ages we are able to constrain better the age of the galaxy, with $R^{2}$ scores of $0.94$.
@@
We show in Fig. 1 the residuals (true property - median posterior distribution) for the test sample as a function of $\log_{10}(M_{*}/\rm{M}_{\odot})$, the mass-weighted age, and $A_{V}$, for the $\tau$-delayed prior. We colour-code the samples with the standard deviation of the posterior distributions for each of the property, showing that simulations with higher residuals have also higher uncertainty in the posterior distributions. We split the data in bins of $\overline{\rm{S/N}}$, including $51<\overline{\rm{S/N}}<$ 5, $105<\overline{\rm{S/N}}<$ 10 and $\overline{\rm{S/N}}>10$. Although there are many more simulations in the last bin, we can see how the performance improves considerably for higher $\overline{\rm{S/N}}$, specially for the mass and dust attenuation. For these three properties we find good agreement between the inferred and simulated properties without significant systematics.
@@
We repeat the analysis in Fig. 2 for the model trained with the Dirichlet prior, finding similar results but with slightly higher scatter for $\log_{10}(M_{*}/\rm{M}_{\odot})$ and $A_{V}$. Our simulations also include very young populations with $\log_{10}\left(t_{50\%}/\mathrm{yr}\right)<8$. These cases correspond to SFHs with very recent bursts, in the last $\sim 100$ Myr of lookback time, and the brightness from young and hot stars prevents from a proper estimation of their previous SFHs, leading to higher residuals and standard deviations in the posteriors for $\log_{10}\left(t_{50\%}/\mathrm{yr}\right)<8$.
@@
We analyse in Appendix A the posterior distributions obtained for the same $\tau$-delayed simulation using the models trained on different priors. We find the expected correlations between parameters, and the widths of the posteriors for the different properties, compared to the widths of the prior distributions, match the $R^{2}$ scores discussed previously. We also perform a simulation-based calibration (SBC) test in Appendix B to verify that the NFs are well trained and thus the posteriors are correctly calibrated.
@@
Figure 1: Residuals of properties compared to the medians of the posterior distributions obtained with the $\tau$-delayed model, for $\log_{10}(M_{*}/\rm{M}_{\odot})$, the $\log_{10}(\rm{age}/{yr})$, referring to the mass-weighted age, and $A_{V}$ [mag]. We split the simulated test sample in bins of mean S/N in the filters F277W, F356W and F [^4] W, 1-5 (left), 5-10 (middle) and $>10$ (right). We colour-code each simulation with the standard deviation of the posterior distribution for the three properties, and include dashed and dotted lines corresponding to the average one and two standard deviations of the posterior distributions respectively. We also plot kernel density distribution contours with black solid lines for clarity.
@@ 5 Inferring properties of JWST galaxies
We select $1083$ galaxies of the HUDF measured in the 19 filters listed in Table 1 and with spectroscopic redshifts. We first examine, across stellar masses and redshifts, the fraction of pixels within two effective radii with a sufficient S/N in order to be fitted with our model. We extract these properties from a catalogue that include stellar masses from surviving stars into a Kron aperture and redshifts from Dahlen et al. (2010); Guo et al. (2012); Oesch et al. (2023) and Bunker et al. (2024), assuming SFHs described by delayed exponentials, with timescales between 1 Myr and 10 Gyr, free metallicity, ages between 1 Myr and the age of the Universe at the redshift of each source, Calzetti attenuation law, and Chabrier IMF.
Again, instead of using the mean or maximum S/N across all available filters, an approach susceptible to dropout effects due to the galaxy redshift, we focus on the mean S/N in the three reddest wide-band filters: F277W, F356W, and F [^4] W. Specifically, we analyse the fraction of pixels with a mean S/N in these three bands equal to or exceeding 5. This threshold, initially selected based on model performance on simulated data, will be further validated with observational data in the following section. Fig. 3 shows that up to redshifts of $z\approx 5-6$ most galaxies with total stellar masses above $10^{8}$ M contain a substantial fraction of pixels above the S/N threshold, allowing them to be analysed effectively with our model.
Figure 3: Distribution of $1.083$ galaxies from the JADES survey in the plane of integrated stellar mass versus spectroscopic redshift, with colours indicating the fraction of pixels within two effective radii (in F [^4] W) that have a mean S/N of 5 or greater in the filters F277W, F356W, and F [^4] W. The lines represent the parameter-space limits used to set thresholds at pixel fractions of 0.20 (red dashed line) and 0.80 (red dotted line), averaged across the full sample. These thresholds indicate, for the ranges of integrated masses and redshifts, the fraction of pixels within the galaxies that can be effectively fitted with our model.
We fit all the pixels of the $1083$ galaxies with the trained model and obtain posterior distributions with $500$ samples for each of them. We analyse statistically these posteriors according to the S/N of the pixels and the redshift in Appendix C. By comparing posterior and prior widths, we find a strong broadening of the posteriors for most of the properties at $\overline{\rm{S/N}}<5$, consistent with the previous results in the synthetic observations, show-casing the regime where our inference is prior-dominated. Moreover, the constraining power for each of the posteriors, measured as the ratio $\sigma_{\rm{posterior}}/\sigma_{\rm{prior}}$, agrees with the $R^{2}$ scores shown in Table 3.
@@ 5.1 Analysis of sample galaxies and comparison with other codes
We select six galaxies from the sample to show the performance of the model across different morphologies, orientations, and redshifts. We include in Table 4 their JADES ID, right ascension, declination, spectroscopic redshift and number of pixels in the segmentation map of the JADES survey with $\overline{\rm{S/N}}>5$.
Table 4: JADES IDs, RA, Dec, spectroscopic redshifts and number of pixels with $\overline{\rm{S/N}}>5$ in the filters F277W, F356W and F [^4] W for the six example galaxies analysed with the model.
@@
Figure 4: Probability distributions for $\log_{10}(M_{*}/\rm{M}_{\odot})$, the mass weighted age [Gyr], and $A_{V}$ [mag] for the pixel with highest mean S/N in the filters F277W, F356W and F [^4] W (in green), and a random pixel out of $1R_{\rm{eff}}$ (in red) for the six galaxies. Step histograms (solid lines) represent the prior distributions (in black) and our posterior distributions (in green and red). The posterior distributions from Prospector are shown with lower opacity (in green and red). The dotted lines correspond to the minimum $\chi^{2}$ parameters obtained with the optimisation code Synthesizer.
@@
In Fig. 4 we show the posterior distributions we obtain for two pixels of each galaxy: one with the highest $\overline{\rm{S/N}}$, and other selected randomly beyond $1R_{\rm{eff}}$, together with the priors, for the model using the $\tau$-delayed prior. We show the surviving stellar mass, the dust attenuation and combine the samples of the posteriors of $\tau$ and $t_{i}$ to derive samples for the mass-weighted age which, as seen in the simulated test sample, is more robust. We omit from this analysis the metallicity, for which we obtain a low $R^{2}$ score and and wide posterior distributions in the simulations.
@@
We compare our posterior distributions with those obtained with Prospector (Johnson et al., 2021b), that are also shown in Fig. 4. We run Prospector with the same assumptions: MILES SSPs, Chabrier IMF and MIST isochrones, $\tau$-delayed SFHs and Calzetti dust attenuation, and the same dust emission, nebular continuum and emission libraries. We also use the priors given in Table 2. We optimise the inference with Dynesty (Speagle, 2020) to reduce the inference time. Despite this optimisation, it takes $\sim 12$ min to fit the posterior distributions with  $2500$ samples. We repeat the inference for the NFs with the same number of samples and it takes  $0.20$ s, which is $\sim 3.5\times 10^{3}$ times faster. This demonstrates the potential of the amortised inference for the pixel-based fitting of high-resolution multi-wavelength observations for large samples of galaxies.
@@
We repeat the fit of all the pixels for the six galaxies included in Table 4 with the optimization code Synthesizer (Pérez-González et al., 2003, 2008). This optimization model employs the BC03 Bruzual & Charlot (2003) stellar population synthesis framework, utilizing the Padova 1994 isochrones (Bertelli et al., 1994) with high-resolution spectra. Both the spectral library and the isochrones differ from those used by our model. However, it assumes the same Chabrier IMF and also incorporates a Calzetti attenuation law and a delayed-$\tau$ model for the SFHs, with the same ranges of parameters than the included in our model, optimizing parameters to minimise the chi-square ($\chi^{2}$) for SED fitting. More information about the nebular emission and continuum models used by Synthesizer can be found in Pérez-González et al. (2003, 2008). We include the parameter estimations of Synthesizer also in Fig. 4 for comparison.
@@
Our posterior distributions are in perfect agreement with those from Prospector, but differ from Synthesizer estimations. The latter is a possible consequence of using a different SPS model. Although we do not have access to the full posterior distributions in Synthesizer, we further analyse how well the models fit the SEDs in the Appendix E. There we perform a posterior predictive check (PPC; Talts et al., 2018), repeating the simulation for every sample of the posteriors for the same two pixels of each of the six galaxies using the parametric model. We compare the results of the PPC with the actual observed photometry, as well as the SED fits from Prospector and Synthesizer, finding an overall good agreement.

### v3

@@ 3.2 Amortised neural inference
We used a neural density estimator known as ’NFs’ to estimate the posterior distributions for the stellar population properties. This method transforms variables described by a simple base distribution, such as a multivariate Gaussian, through a series of parametrised invertible transformations. The parameters of these transformations are trained to minimise the Kullback–Leibler divergence between the target distribution and the model’s estimate. For a more detailed explanation, we refer to Jimenez Rezende & Mohamed (2015) and Hahn & Melchior (2022).
We used a specific implementation of NFs known as MAF (masked autoregressive flow; Papamakarios et al., 2017), implemented with the SBI module (Tejero-Cantero et al., 2020). Following the example of Hahn & Melchior (2022), we selected 15 MADE blocks (masked autoencoder for distribution estimation), each containing two hidden layers with 500 units. We split the data into training and validation sets in a 90/10 ratio and used the Adam optimiser (Kingma & Ba, 2014) with a learning rate of $5\cdot 10^{-4}$. To prevent overfitting, we monitored the model’s performance on the validation set and halted the training if the validation likelihood did not improve after $20$ epochs.
We performed two trainings for the two different simulations, with parametric (a) and non-parametric (b) SFHs. The output of the network are the posterior distributions for the following properties: the formed stellar mass, surviving stellar mass including remnants, average SFR for the last $100$ Myr, metallicity $[M/H],$ and dust attenuation index, $A_{V}$. For the SFHs, in (a) we inferred $\tau$ and $t_{i}$, both in Gyr, while in (b) we inferred the normalised time $(t_{x}/t_{\rm{universe(z)}})$ at which $25\%$, $50\%,$ and $75\%$ of the total stellar mass were formed.
We discuss the possibility of also retrieving the redshift pixel-by-pixel, as done in (Pérez-González et al., 2023) in Sect.8. However, in this work we introduce the aforementioned integrated spectroscopic redshifts as an input in the network together with the photometry. The training was performed with $1.000.000$ simulated CSPs. We generated another $10.000$ for testing purposes.
@@ 3.3 Fitting of JADES galaxies
We use the trained model to fit the galaxies introduced in Sect. 2. We computed the noise of the galaxies in each pixel and filter as $\sqrt{\sigma^{2}+\sigma_{b}^{2}}$, where $\sigma$ is the flux uncertainty from the mosaics and $\sigma_{b}$ the standard deviation of the background signal. We refer to the NIRCam IDs of the galaxies throughout the paper. All spectroscopic redshifts are extracted from catalogues (Dahlen et al., 2010; Guo et al., 2012; Oesch et al., 2023; Bunker et al., 2024). We input the photometry with uncertainties (in AB magnitudes) into the network, again repeating the cut at a $1\sigma$ depth limit for each filter as done for the simulations, as well as the redshifts. Thus, we obtained the posterior distributions for each pixel for (a) $\log_{10}(M_{*}/\rm{M}_{\odot})$, $\log_{10}(M_{*}^{\rm{formed}}/\rm{M}_{\odot})$, $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)), $\tau$, $t_{i}$, [MM/H] and $A_{V}$; and (b) $\log_{10}(M_{*}/\rm{M}_{\odot})$, $\log_{10}(M_{*}^{\rm{formed}}/\rm{M}_{\odot})$, $\log_{10}$(SFR/($\rm{M}_{\odot}$ yr-1)), $t_{25\%}$, $t_{50\%}$, $t_{75\%}$, [MM/H] and $A_{V}$.
The results are organised as follows. In Sect. 4, we analyse the performance of the model in the simulated test set. In Sect. 5, we describe our fit of the individual pixels from galaxies of the JADES survey. We inspect the posterior distributions at different redshifts (Sect. 5.2), comparing them with other SED fitting codes. Then, in Sect. 6, we present the maps for stellar population properties of selected galaxies. Finally, in Sect. 6.2, we describe our statistical study and compare the integrated stellar masses with their pixel-based analogues to investigate the outshining phenomenon.
@@ 4 Recovering properties of the simulated stellar populations
We trained the neural density estimator over $\sim 1$ hour and $30$ min. We obtained the posterior probability distributions for each stellar population parameter for the full test set, from the photometry with the associated uncertainties. It takes $\sim 0.04$ s to obtain $500$ samples of the distributions for all the properties for one simulation. All of the time computations were performed on a single core of an Apple M3 Max CPU, part of an ARM-based 14-core system with 36 GB of unified memory. However, this model natively supports GPU execution, which can further reduce inference time.
@@
To assess the accuracy of the model, we computed the medians of the posteriors inferred for the parameters and compared them with the true simulated values for the full test sample with S/N averaged in the filters F277W, F356W, and F [^4] W (hereafter, $\overline{\rm{S/N}}$) higher than 1. This ensures we did not show any simulations that are completely above of the observational capabilities of ACS and NIRCam, with upper limits for most of the filters. In Table 3, we include the $R^{2}$ values [^3] Also known as coefficient of determination $R^{2}=1-\frac{\sum_{i=1}^{n}\left(\theta_{i}-\widehat{\theta}_{i}\right)^{2}}{\sum_{i=1}^{n}\left(\theta_{i}-\overline{\theta}_{i}\right)^{2}}$, not to be confused with the square of the Pearson’s coefficient. for the training using parametric (a) and non-parametric (b) SFHs. The performance is relatively good for the stellar mass and the dust attenuation index, but worse for the SFR, metallicity, and ages as a result of the degeneracies that we know exist being reproduced when inferring these properties from photometry. As expected, the $R^{2}$ values for the SFR and [MM/H] are higher for parametric SFHs. However, while $t_{i}$ can be well constrained, the accuracy of the medians of the posterior distribution for $\tau$ are extremely low. This can be explained not only because it is difficult to retrieve from photometry, but also because it is very degenerated with $t_{i}$ (Conroy, 2013). Instead, we find that by combining $\tau$ and $t_{i}$ samples to obtain mass-weighted ages, we are able to constrain better the age of the galaxy, with $R^{2}$ scores of $0.94$.
@@
We show in Fig. 1 the residuals (true property: median posterior distribution) for the test sample as a function of $\log_{10}(M_{*}/\rm{M}_{\odot})$, the mass-weighted age, and $A_{V}$, for the $\tau$-delayed prior. We colour-coded the samples with the standard deviation of the posterior distributions for each of the property, showing that simulations with higher residuals have also higher uncertainty in the posterior distributions. We split the data into bins of $\overline{\rm{S/N}}$, including $1<\overline{\rm{S/N}}<5$, $5<\overline{\rm{S/N}}<10$, and $\overline{\rm{S/N}}>10$. Although there are many more simulations in the last bin, we can see how the performance improves considerably for higher $\overline{\rm{S/N}}$, especially for the mass and dust attenuation. For these three properties, we find a good agreement between the inferred and simulated properties without significant systematics, except for a mild apparent trend in $A_{V}$. The latter reflects the influence of the prior shape under low S/N conditions, with the posteriors clustering around the midpoint of the uniform prior range.
@@
We repeat the analysis in Fig. 2 for the model trained with the Dirichlet prior, finding similar results, but with slightly higher scatter for $\log_{10}(M_{*}/\rm{M}_{\odot})$ and $A_{V}$. Our simulations also include very young populations with $\log_{10}\left(t_{50\%}/\mathrm{yr}\right)<8$. These cases correspond to SFHs with very recent bursts, in the last $\sim 100$ Myr of lookback time; furthermore, the brightness from young and hot stars prevents us from making a proper estimation of their previous SFHs, leading to higher residuals and standard deviations in the posteriors for $\log_{10}\left(t_{50\%}/\mathrm{yr}\right)<8$.
The galaxies exhibiting larger scatter typically show degeneracies between stellar mass, age, and dust attenuation: older, dustier stellar populations can resemble younger, but less massive, less extinguished ones in their observed SEDs. Conversely, very young, dust-free populations can mimic older and dustier systems with higher stellar masses. This degeneracy often leads to bimodal posterior distributions, where the data cannot distinguish between two distinct combinations of physical parameters. Instead of collapsing this ambiguity into a single best-fit estimate (or the median, as done in Fig. 2) that is not representative, our model returns the full posterior. This captures the complete range of plausible solutions, including bimodal ones, allowing us to directly characterise the associated uncertainties
@@
We analyse in Appendix A the posterior distributions obtained for the same $\tau$-delayed simulation using the models trained on different priors. We find the expected correlations between parameters and the widths of the posteriors for the different properties (compared to the widths of the prior distributions) match the $R^{2}$ scores discussed previously. We also performed a simulation-based calibration (SBC) test in Appendix B to verify that the NFs are well trained and, thus, the posteriors are correctly calibrated.
@@
Figure 1: Residuals of properties compared to the medians of the posterior distributions obtained with the $\tau$-delayed model, for $\log_{10}(M_{*}/\rm{M}_{\odot})$, the $\log_{10}(\rm{age}/{yr})$, referring to the mass-weighted age, and $A_{V}$ [mag]. We split the simulated test sample in bins of mean S/N in the filters F277W, F356W, and F [^4] W, 1-5 (left), 5-10 (middle) and >10>10 (right). We colour-coded each simulation with the standard deviation of the posterior distribution for the three properties and included dashed and dotted lines corresponding to the average one and two standard deviations of the posterior distributions respectively. We also plotted kernel density distribution contours with black solid lines for clarity.
@@ 5 Inferring properties of JWST galaxies
5.1 Sample selection
We selected $1083$ galaxies of the HUDF measured in the 19 filters listed in Table 1 and with spectroscopic redshifts. We first examined, across stellar masses and redshifts, the fraction of pixels within two effective radii with a sufficient S/N in order to be fitted with our model. We extracted these properties from a catalogue that includes stellar masses from surviving stars into a Kron aperture and redshifts from Dahlen et al. (2010); Guo et al. (2012); Oesch et al. (2023), and Bunker et al. (2024), assuming SFHs described by delayed exponentials, with timescales between 1 Myr and 10 Gyr, free metallicity, ages between 1 Myr and the age of the Universe at the redshift of each source, the Calzetti attenuation law, and a Chabrier IMF.
Again, instead of using the mean or maximum S/N across all available filters, an approach susceptible to dropout effects due to the galaxy redshift, we focus on the mean S/N in the three reddest wide-band filters: F277W, F356W, and F [^4] W. Specifically, we analysed the fraction of pixels with a mean S/N in these three bands equal to or exceeding 5. This threshold, initially selected based on model performance on simulated data, will be further validated with observational data in the following section. Fig. 3 shows that up to redshifts of $z\approx 5-6$ most galaxies with total stellar masses above $10^{8}$ M contain a substantial fraction of pixels above the S/N threshold, allowing them to be analysed effectively with our model.
Figure 3: Distribution of $1.083$ galaxies from the JADES survey in the plane of integrated stellar mass versus spectroscopic redshift, with colours indicating the fraction of pixels within two effective radii (in F [^4] W) that have a mean S/N of 5 or greater in the filters F277W, F356W, and F [^4] W. The lines represent the parameter-space limits used to set thresholds at pixel fractions of 0.20 (red dashed line) and 0.80 (red dotted line) averaged across the full sample. These thresholds indicate, for the ranges of integrated masses and redshifts, the fraction of pixels within the galaxies that can be effectively fitted with our model.
We fit all the pixels of the $1083$ galaxies with the trained model and obtained posterior distributions with $500$ samples for each of them. We carried out a statistical analysis of these posteriors according to the S/N of the pixels and the redshift in Appendix C. By comparing the posterior and prior widths, we found a strong broadening of the posteriors for most of the properties at $\overline{\rm{S/N}}<5$, consistent with the previous results in the synthetic observations. This showcases the regime where our inference is prior-dominated, namely, where the observational data provide limited constraining power and the resulting posterior distribution closely follows the form of the adopted prior. Moreover, the constraining power for each of the posteriors, measured as the ratio $\sigma_{\rm{posterior}}/\sigma_{\rm{prior}}$, agrees with the $R^{2}$ scores given in Table 3.
@@ 5.2 Analysis of sample galaxies and comparison with other codes
We selected six galaxies from the sample to show the performance of the model across different morphologies, orientations, and redshifts. In Table 4, we include their JADES ID, right ascension, declination, spectroscopic redshift, and number of pixels in the segmentation map of the JADES survey with $\overline{\rm{S/N}}>5$.
Table 4: JADES IDs, RA, Dec, spectroscopic redshifts, and number of pixels with $\overline{\rm{S/N}}>5$ in the filters F277W, F356W, and F [^4] W for the six example galaxies analysed with the model.
@@
Figure 4: Probability distributions for $\log_{10}(M_{*}/\rm{M}_{\odot})$, the mass-weighted age [Gyr], and $A_{V}$ [mag] for the pixel with highest mean S/N in the filters F277W, F356W, and F [^4] W (in green), and a random pixel out of $1R_{\rm{eff}}$ (in red) for the six galaxies. Step histograms (solid lines) represent the prior distributions (in black) and our posterior distributions (in green and red). The posterior distributions from Prospector are shown with lower opacity (in green and red). The dotted lines correspond to the minimum $\chi^{2}$ parameters obtained with the optimisation code Synthesizer.
@@
In Fig. 4, we show the posterior distributions we obtain for two pixels of each galaxy: one with the highest $\overline{\rm{S/N}}$ and other selected randomly beyond $1R_{\rm{eff}}$, together with the priors, for the model using the $\tau$-delayed prior. We show the surviving stellar mass and the dust attenuation, combining the samples of the posteriors of $\tau$ and $t_{i}$ to derive samples for the mass-weighted age which, as seen in the simulated test sample, is more robust. We omitted from this analysis the metallicity, for which we obtained a low $R^{2}$ score and and wide posterior distributions in the simulations.
@@
We compared our posterior distributions with those obtained with Prospector (Johnson et al., 2021b), which are also shown in Fig. 4. We ran Prospector with the same assumptions: MILES SSPs, Chabrier IMF and MIST isochrones, $\tau$-delayed SFHs and Calzetti dust attenuation, and the same dust emission, nebular continuum, and emission libraries. We also used the priors given in Table 2. We optimised the inference with Dynesty (Speagle, 2020) to reduce the inference time. Despite this optimisation, it takes $\sim 12$ min to fit the posterior distributions with $2500$ samples. We repeated the inference for the NFs with the same number of samples and it takes $0.20$ s, which is $\sim 3.5\times 10^{3}$ times faster. This demonstrates the potential of the amortised inference for the pixel-based fitting of high-resolution multi-wavelength observations for large samples of galaxies.
@@
We repeated the fit of all the pixels for the six galaxies included in Table 4 with the optimisation code Synthesizer (Pérez-González et al., 2003, 2008). This optimisation model employs the BC03 Bruzual & Charlot (2003) stellar population synthesis framework, utilising the Padova 1994 isochrones (Bertelli et al., 1994) with high-resolution spectra. Both the spectral library and the isochrones differ from those used by our model. However, it assumes the same Chabrier IMF and also incorporates a Calzetti attenuation law and a delayed-$\tau$ model for the SFHs, with the same ranges of parameters than the included in our model, optimising the parameters to minimise the chi-square ($\chi^{2}$) for SED fitting. More information about the nebular emission and continuum models used by Synthesizer can be found in Pérez-González et al. (2003, 2008). We also give the parameter estimations of Synthesizer in Fig. 4 for comparison.
@@
Our posterior distributions are in perfect agreement with those from Prospector, but differ from Synthesizer estimations. The latter is a possible consequence of using a different SPS model. Although we do not have access to the full posterior distributions in Synthesizer, we further analysed how well the models fit the SEDs in the Appendix E. There, we performed a posterior predictive check (PPC; Talts et al., 2018), repeating the simulation for every sample of the posteriors for the same two pixels of each of the six galaxies using the parametric model. We compared the results of the PPC with the actual observed photometry, as well as the SED fits from Prospector and Synthesizer, finding an overall good agreement.